The molecular weight of MgCl2 is 95. How many milligrams is 1mEq per liter? The molecular...

  1. The molecular weight of MgCl2 is 95. How many milligrams is 1mEq per liter?
  2. The molecular weight of NaH2PO4.H2O is 138. How many milligrams is 1mEq per liter?
Answer #1

Answer: 1 mEq means ( 1 milli equivalent) means 1/1000 times the equivalent mass of given compound.

Molar mass of MgCl2 = 95 g/mol (given)

Since Mg is di-positive cation (Mg+2) so the equivalent mass of MgCl2 = 95/2 g = 47.5 g/equivalent

1 mEq of MgCl2 = (1/1000) x 47.5 g = 47.5 x 10-3 g = 47.5 mg

Molar mass of NaH2PO4.H2O = 138 g/mol (given)

Since Na is mono-positive cation (Na+) so the equivalent mass of NaH2PO4.H2O = 138/1 g = 138 g/equivalent

1 mEq of NaH2PO4.H2O = (1/1000) x 138 g = 138 x 10-3 g = 138 mg